OpenStudy (anonymous):

how do i find the product of -3(x+6) ?

OpenStudy (amistre64):

same as any other multiplication, but lets take a simpler example: 3(8) = 24 right?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

You distribute the -3 which means you multiply the -3 times each term in the parentheses. -3 * x = -3x and -3 * 6 = -18 so the answer is -3x - 18

OpenStudy (amistre64):

then: 3(6+2) should equal 24 right?

OpenStudy (anonymous):

right...

OpenStudy (amistre64):

good, then lets "distribute" the 3 into the (6+2) and see if it works 3(6) + 3(2) = 24 18 + 6 = 24 24 = 24 Does that make sense to you?

OpenStudy (anonymous):

o simple...yes it does...thanks !

OpenStudy (amistre64):

youre welcome... now just do that to your original problem and you get your answer :)

OpenStudy (amistre64):

-3(x+6) -3(x) + -3(6) -3x + -18 -3x - 18

OpenStudy (anonymous):

then do i divide -3x by -3 abd -18/-3 ?

OpenStudy (amistre64):

if you want to "factor" the -3 back out again, then yes.... divide at all by -3

OpenStudy (amistre64):

wait; your not trying to solve for x in this equation

OpenStudy (amistre64):

simply put, it aint an equation...there is no (=) sign to equate anything to..

OpenStudy (amistre64):

-3(x+6) is just an empty term waiting for something... -3(x+6) = 9 is something that we can figure out because it gives us the ability to "equate" one side with another

OpenStudy (amistre64):

make sense?

OpenStudy (anonymous):

ya that does...Thank you so much!